### v2

# Cross-Modal Consistency Learning for Sign Language Recognition

Abstract
Pre-training has been proven to be effective in boosting the performance of Isolated Sign Language Recognition (ISLR). Existing pre-training methods solely focus on the compact pose data, which eliminates background perturbation but inevitably suffers from insufficient semantic cues compared to raw RGB videos. Nevertheless, learning representation directly from RGB videos remains challenging due to the presence of sign-irrelevant visual features. To address this dilemma, we propose a Cross-modal Consistency Learning framework (CCL-SLR), which leverages cross-modal consistency between both RGB and pose modalities in a self-supervised paradigm. First, CCL-SLR employs contrastive learning for instance discrimination within and across modalities. Through single-modal and cross-modal contrastive learning, CCL-SLR gradually aligns the feature spaces of RGB and pose modalities, thereby extracting consistency sign representation. Second, we further introduce Motion-Preserving Masking (MPM) and Semantic Positive Mining (SPM) techniques to improve cross-modal consistency from the perspective of data augmentation and sample clustering, respectively. Extensive experiments on four ISLR benchmarks show that CCL-SLR achieves impressive performance, demonstrating its effectiveness. The code will be released to the public.
1 Introduction
Sign language, which combines body gestures and facial expressions to convey semantic meaning, serves as the primary communication tool within the Deaf community. To facilitate a barrier-free understanding between sign language and spoken language, extensive research [17, 44, 28, 21, 55, 13, 16, 46, 48] has explored sign language understanding from various perspectives. Isolated Sign Language Recognition (ISLR) is a fundamental task in sign language understanding, that focuses on recognizing the meaning of sign language videos at the word level to achieve fine-grained understanding. For the deployment of ISLR models in real-world applications, developing robust visual encoders with enhanced representation capabilities has increasingly been recognized as a crucial research direction.
Figure 1: Comparison between previous pre-training paradigms and ours. (a) Previous paradigms [46, 48, 13, 16] ignore modeling the interaction between RGB and pose modalities during pre-training. (b) Our proposed method effectively aligns embedding spaces of two modalities via cross-modal consistency learning.
Current ISLR methods mainly employ supervised paradigms on small-scale datasets [26, 24, 18, 15], which can be categorized into two paradigms based on the input modalities. (1) RGB-based methods [28, 26, 27, 25, 24, 14, 15, 29] explore various strategies to extract discriminative visual features from RGB videos. Although these methods are intuitive and straightforward, they fail to explicitly capture key gesture-related information. Furthermore, signer-dependent variations hinder effective feature extraction, ultimately compromising the robustness of the sign representations. (2) Pose-assisted methods [17, 44, 21, 20, 55] integrate the pose modality and employ multi-stream fusion frameworks to leverage cross-modal complementary information. By utilizing the compact pose modality, these methods efficiently eliminate non-skeletal noise. However, current pose data is directly derived from general-domain pose extractors. Due to motion blur and hand occlusion in sign language videos, these general-domain extractors inevitably produce errors and missing detections, limiting the performance ceiling of these pose-assisted methods.
Although these supervised methods have achieved promising results, the scarcity of labeled training data inherently constrains their representation and generalization capabilities. To tackle this problem, recent works [13, 16, 46, 48, 47] attempt to explore self-supervised pre-training strategies in the pose modality to enhance visual modeling capabilities. For instance, MASA [47] proposes a motion-aware masked autoencoder to utilize the pose modality by introducing a motion prior and global semantic cues into pre-training. However, it incorporates only RGB information through late fusion during fine-tuning, failing to leverage original RGB features during pre-training. As shown in Fig. 1(a), these self-supervised methods focus solely on extracting visual information from pose modality. This paradigm fails to consider cross-modal feature interactions and overlooks the potential benefits of original RGB visual information.
In this paper, we aim to introduce cross-modal consistency learning into the pre-training paradigm to fully enhance feature extraction in subsequent stages. Notably, effectively applying self-supervised paradigms such as contrastive learning to cross-modal interaction is non-trivial due to two key challenges: (1) Redundant information in the RGB modality. RGB images contain redundant non-semantic information in background regions, such as signer-dependent variations such as body type and skin color, which are often irrelevant to pose information. These redundant cues disrupt the extraction of sign language-related features, further impacting the ISLR performance. (2) Inconsistency between modalities. Although the RGB and pose data of the same instance correspond to each other, the similarity of structural relationships between instances across the two modalities are inconsistent, leading to the inherent misalignment in feature distributions across RGB and pose modalities.
To this end, we propose a novel Cross-modal Consistency Learning framework (CCL-SLR), which aims to model consistency constraints during pre-training, as shown in Fig. 1 (b). We employ single-modal contrastive learning and cross-modal contrastive learning to facilitate instance discrimination within individual modalities and across different modalities, respectively. To more effectively extract consistent sign language representations from both RGB and pose modalities in the contrastive learning paradigm, our approach incorporates two key modules: Motion-Preserving Masking (MPM) and Semantic Positive Mining (SPM). (1) To eliminate impact from redundant features, MPM introduces an off-the-shelf flow-based model [31] to extract latent representations, leveraging this information to capture motion dynamics. By preserving motion-related information, MPM suppresses non-semantic regions in RGB videos, thereby enhancing the relevance of learned features. (2) To capture cross-modal structural consistency, SPM identifies semantically similar samples within each modality to provide pseudo-supervision for both modalities. Through a multi-label classification approach, this pseudo-supervision introduces shared positive constraints across different modalities, effectively bridging the distribution gap between them. To validate the effectiveness of our method, we conduct extensive experiments on four widely adopted ISLR benchmarks. In summary, our contributions are as follows:
  • We propose a novel pre-training framework named CCL-SLR, which effectively leverages the single-modal and cross-modal contrastive learning on both RGB and pose modalities for enhancing ISLR.
  • We introduce MPM and SPM, two innovative techniques designed to mitigate the distribution gap between modalities and further strengthen the consistency constraints during cross-modal pre-training.
  • We conduct extensive experiments on four widely-adopted benchmarks and achieve impressive performance, which demonstrates effectiveness and representation capability of CCL-SLR.
2 Related Work
Figure 2: (a) The CCL-SLR pre-training framework. RGB and pose sequences are processed through two single-modal branches to learn modality-specific representation. During this process, MPM is applied to mask motion-irrelevant regions in the RGB videos. Additionally, cross-modal positive samples are identified through SPM, which are utilized for single-modal loss $\mathcal{L}^{R}$ and $\mathcal{L}^{L}$. Finally, the consistency of RGB and pose embedding spaces is constrained through cross-modal contrastive loss $\mathcal{L}_{cross}$. (b) An overview of the Semantic Positive Mining. SPM selects the top-$k$ most similar samples from each modality-specific memory bank and merges them to get pseudo-labels.
Sign Language Recognition. Sign language recognition is a subtask in the field of sign language understanding, which aims to recognize a single sign word within a given video. Early works [6, 7, 37, 32, 34] utilize hand-crafted features to model spatio-temporal representation. With the development of deep convolution neural networks (CNNs), numerous recent works [26, 27, 25, 24, 14, 15, 29] leverage CNN-based architectures such as TSM [30], R3D [10], and I3D [1] to extract visual representations from RGB videos. For example, StepNet [36] uses TSM to model part-level spatial and temporal relationships, achieving impressive performance on ISLR.
However, these RGB-based methods may suffer from static irrelevant cues in sign language videos [55], such as background and clothing. Consequently, several works have begun exploring pose-assisted methods, leveraging the compact and informative nature of sign pose data. For instance, NLA-SLR [55] presents a video-keypoint multi-stream backbone and utilize semantic information within sign glosses to promote ISLR. SAM-v2 [21] jointly leverages pre-extracted multi-modal data with a branching ensemble model to achieve a higher recognition rate. Although these works effectively fuse RGB and pose modality-specific features, they rely on complex model designs and feature fusion strategies, which limits their scalability. Meanwhile, some studies [13, 16, 47, 46, 48] incorporate self-supervised pre-training paradigms into ISLR to address the challenge of data scarcity. SignBERT [13] enhances model representation capacity through a generative pretext task specifically designed for pose data during self-supervised pre-training. BEST [46] leverage discrete variational autoencoder to compress hand shapes into a codebook for implicit clustering, serving as pseudo-labels to better learn skeletal priors. Nevertheless, these approaches overlook the benefits of aligning consistency visual representations of RGB and pose data during pre-training for downstream ISLR tasks.
In this work, we focus on the ISLR pre-training paradigm, learning instance discriminative representations through contrastive learning within and across modalities.
Contrastive Representation Learning. Contrastive learning has significantly advanced self-supervised visual representation learning by performing instance discrimination in a fully self-supervised manner [11, 3, 2, 43, 8, 33]. It pulls positive samples closer and pushes negative samples away. Since there are no supervised labels, different data augmentations of the same instance are treated as a positive pair, while samples from other instances are used as negative pairs. Recently, MoCo [11] and MoCo v2 [3] have strengthened contrastive learning by maintaining a dynamic memory bank of negative sample embedding from previous iterations. Building on these methods, current advances have demonstrated substantial improvements in representation capability through various technical innovations, including positive mining [54, 52, 42, 19] and data augmentation [45, 5, 38, 53].
Despite their success in various downstream tasks, integrating the contrastive learning paradigm into multi-modal ISLR pre-training remains challenging. To this end, we propose CCL-SLR, which extends the MoCo v2 framework to the challenging task of multi-modal sign language representation learning.
3 Method
As illustrated in Fig. 2, our framework first applies modality-specific data augmentation to generate augmented samples. Subsequently, these augmented samples are fed into their respective encoders for feature extraction. The encoder outputs are then utilized to compute the contrastive loss, facilitating the learning of instance discriminative representations within and across modalities. For the RGB modality, we introduce a novel Motion-Preserving Masking technique to suppress redundant non-semantic visual cues in RGB videos. Furthermore, to enhance single-modal representation learning while leveraging cross-modal consistency, we propose a Semantic Positive Mining strategy that identifies semantically similar positive pairs.
3.1 Preliminaries
Since CCL-SLR involves two modalities, we first introduce contrastive learning in the single-modal case to simplify the explanation. Single-modal contrastive learning has emerged as a powerful paradigm across various tasks, with notable success demonstrated by MoCo v2 [3]. Given a single-modal input $x$ in a mini-batch, the learning process begins with data augmentation to generate two distinct samples $x_{q}$ and $x_{k}$ (query and key). Two encoders, $\mathcal{E}_{q}$ and $\mathcal{E}_{k}$, are then applied to embed $x_{q}$ and $x_{k}$ into a hidden space:
$z_{q}=\mathcal{E}_{q}(x_{q};\theta_{q}),\quad z_{k}=\mathcal{E}_{k}(x_{k};\theta_{k}),$ (1)
where $\mathcal{E}_{q}$ and $\mathcal{E}_{k}$ denote the query and key encoders, $\theta_{q}$ and $\theta_{k}$ denote the learnable parameters of the encoders. Notably, the key encoder $\mathcal{E}_{k}$ is maintained as an exponential moving average (EMA) of the query encoder $\mathcal{E}_{q}$, updated as:
$\theta_{k}\leftarrow\gamma\theta_{k}+(1-\gamma)\theta_{q},$ (2)
where $\gamma$ denotes the momentum coefficient. During the pre-training phase, instance discrimination is achieved through the InfoNCE [39] loss:
$\small\mathcal{L}_{cl}(z_{q},z_{k},\mathcal{S})=-\log\frac{\exp(z_{q}{\cdot}z_{k}/\tau)}{\exp(z_{q}{\cdot}z_{k}/\tau)+\sum\limits_{s_{i}\in\mathcal{S}}\exp(z_{q}{\cdot}s_{i}/\tau)},$ (3)
where $\tau$ represents the temperature coefficient [12], and $\mathcal{S}$ denotes the negative sample embedding set maintained in a memory bank of size $N$. The similarity between samples is computed using the dot product operation.
3.2 Single-Modal Contrastive Learning
Framework Architecture. As illustrated in Fig. 2, CCL-SLR consists of two parallel single-modal branches: an RGB branch and a pose branch. Each branch encompasses two pathways for key and query samples, along with two queue-based memory banks. Each pathway consists of an encoder and a projection head. We extract the corresponding pose sequence $X^{P}\in\mathbb{R}^{T\times J\times C_{p}}$ from the RGB frame sequence $X^{R}\in\mathbb{R}^{T\times H\times W\times 3}$ using an off-the-shelf estimator [4]. Here, $T$ denotes the frame length, $J$ denotes the number of joints, and $H$ and $W$ correspond to the spatial dimensions. $C_{p}=3$ represents the XY coordinates along with the confidence score of pose data.
First, we use spatial-temporal augmentation to generate different query and key samples for each single-modal branch, denoted as $X^{m}_{s}$, where $m\in\{R,P\}$ represents RGB and pose modalities and $s\in\{q,k\}$ represents query and key samples. Second, each sample $X^{m}_{s}$ passes through its corresponding encoder $\mathcal{E}^{m}_{s}$ to obtain an embedding ${g}^{m}_{s}\in\mathbb{R}^{d}$. The RGB encoder employs R3D-50 [10] backbone, while the pose encoder utilizes a GCN+Transformer architecture [50]. Then, ${g}^{m}_{s}$ is projected using a multi-layer perceptron (MLP) projection head $\mathcal{P}^{m}_{s}$ to obtain the final embedding $f^{m}_{s}\in\mathbb{R}^{d}$. Two queue-based memory banks $\mathcal{M}^{m}_{g}\in\mathbb{R}^{N\times d}$ and $\mathcal{M}^{m}_{f}\in\mathbb{R}^{N\times d}$ are utilized for storing L2-normalized embeddings of ${g}^{m}_{k}$ and ${f}^{m}_{k}$, respectively:
$\|{g}^{m}_{k}\|_{2}\overset{Enqueue}{\Longrightarrow}\mathcal{M}^{m}_{g},\quad\|{f}^{m}_{k}\|_{2}\overset{Enqueue}{\Longrightarrow}\mathcal{M}^{m}_{f},$ (4)
where $\|\cdot\|_{2}$ denotes the L2 normalization operation. Here, $\mathcal{M}^{m}_{g}$ is utilized to mine semantic positive samples, and $\mathcal{M}^{m}_{f}$ is applied for contrastive learning.
Figure 3: An overview of MPM. $h(\cdot)$ and $h^{-1}(\cdot)$ denote the transformation from pixel/latent space to latent/pixel space, respectively. The video is encoded into latent space using a flow-based generative model. Channels in the latent representation with low standard deviation along the temporal dimension are then masked out. The masked latent representation is decoded back into pixel space. The reconstructed frames are then processed through gray-scaling and binarization to produce motion-aware mask sequence.
Motion-Preserving Masking (MPM). Unlike pose sequences, which focus on keypoint coordinates, RGB videos contain substantial non-gestural information such as background and clothing. To preserve essential motion components while suppressing irrelevant regions, we propose MPM, a data augmentation technique that enhances motion capture capabilities and mitigates bias from non-skeletal elements. As shown in Fig.3, MPM employs a pre-trained flow-based generative model, AdvFlow [31], to transform original RGB videos into motion-preserving videos and generate a mask sequence for data augmentation.
Concretely, given an RGB video $X^{R}$, we encode it into a latent representation $Z\in\mathbb{R}^{T\times d_{z}}$. Empirically, the temporal standard deviation of each channel reliably indicates its motion content. We leverage this measure on $Z$ to preserve dynamic features while suppressing static ones:
$\hat{Z}=Z\otimes\mathds{1}(\sigma(Z)<p),$ (5)
where $\mathds{1}(\cdot)$ denotes the indicator function, $\otimes$ denotes the element-wise multiplication, $\sigma(\cdot)$ computes the standard deviation along the temporal channel, and $p$ controls the mask intensity. The motion-preserving video $\hat{X}^{R}\in\mathbb{R}^{T\times H\times W\times 3}$ is then decoded from the masked latent representation $\hat{Z}$. Subsequently, we derive the binary mask sequence $\zeta\in\mathbb{R}^{T\times H\times W}$ from $\hat{X}^{R}$ and apply it to $X^{R}$ with probability $\alpha$ to generate the augmented sample.
Semantic Positive Mining (SPM). While single-modal contrastive learning enables CCL-SLR to extract modality-specific representations, the “semantic collision” problem identified by CrosSCLR [54] remains a challenge. The semantic collision occurs during contrastive learning when sample embeddings with high semantic similarity in the memory bank are mistakenly treated as negative examples. To mitigate this semantic misalignment, we introduce SPM, a mechanism that identifies semantic positive samples across memory banks of different modalities. As illustrated in Fig. 2 (b), we select the top-$k$ nearest neighbors of ${g}^{m}_{q}$ from $\mathcal{M}^{m}_{g}$ as positive samples, while treating the remaining embeddings as negative samples. This selection process generates binary pseudo-labels $y^{m}\in\mathbb{R}^{N}$ for both modalities:
$y^{R}=Top\text{-}k({g}^{R}_{q}\cdot\mathcal{M}^{R}_{g}),\quad y^{P}=Top\text{-}k({g}^{P}_{q}\cdot\mathcal{M}^{P}_{g}).$ (6)
To align the neighborhood structure across modalities, we combine $y^{R}$ and $y^{P}$ by taking their element-wise maximum to obtain the cross-modal binary pseudo-label $y$. The embeddings in $\mathcal{M}^{m}_{f}$ serve as classifiers, generating logits $\hat{y}^{m}\in\mathbb{R}^{N}$ that classify ${f}^{m}_{q}$ into $N$ pseudo categories:
$\hat{y}^{m}=\phi(f^{m}_{q}\cdot\mathcal{M}_{f}^{m}/\tau_{spm}),$ (7)
where $\tau_{spm}$ denotes the temperature coefficient and $\phi(\cdot)$ is the sigmoid function [51]. Then, $\hat{y}^{m}$ and $y$ are utilized for binary cross-entropy (BCE) loss to help single-modal contrastive learning bring the query sample closer to semantic positive samples from both modalities:
$\small\mathcal{L}^{m}_{spm}=-\frac{1}{N}\sum_{i=1}^{N}\left[y_{i}\log\hat{y}^{m}_{i}+(1-y_{i})\log(1-\hat{y}^{m}_{i})\right],$ (8)
Additionally, to perform intra-modal instance discrimination, we formulate the single-modal contrastive loss for both modalities as follows:
$\mathcal{L}^{m}_{cl}=\mathcal{L}_{cl}({f}^{m}_{q},{f}^{m}_{k},\mathcal{M}^{m}_{f}).$ (9)
The final single-modal loss is formulated as::
$\mathcal{L}_{single}=\underbrace{\mathcal{L}^{R}_{cl}+\mathcal{L}^{R}_{spm}}_{\displaystyle\mathcal{L}^{R}}+\underbrace{\mathcal{L}^{P}_{cl}+\mathcal{L}^{P}_{spm}}_{\displaystyle\mathcal{L}^{P}}.$ (10)
3.3 Cross-Modal Contrastive Learning
While previous work [49] successfully employed Kullback-Leibler (KL) divergence to ensure distribution consistency within the pose modality, cross-modal alignment between pose and RGB modality remains challenging due to the substantial distribution gap between them. To this end, we extend single-modal contrastive learning to the cross-modal domain by directly aligning embeddings from both modalities in the latent space. Specifically, for RGB features $f^{R}_{q}$, we utilize $f^{P}_{k}$ as positive samples while treating embeddings in $\mathcal{M}^{P}_{f}$ as negative. A symmetric procedure is applied to pose features. The bidirectional cross-modal contrastive loss functions are defined as:
$\mathcal{L}^{R\rightarrow P}$ $=\mathcal{L}_{cl}(f^{P}_{q},f^{R}_{k},\mathcal{M}^{R}_{f}),$ (11)
$\mathcal{L}^{P\rightarrow R}$ $=\mathcal{L}_{cl}(f^{R}_{q},f^{P}_{k},\mathcal{M}^{P}_{f}),$ (12)
$\mathcal{L}_{cross}$ $=\mathcal{L}^{R\rightarrow P}+\mathcal{L}^{P\rightarrow R}.$ (13)
The overall pre-training objective function is formulated as:
$\mathcal{L}=\mathcal{L}_{single}+\mathcal{L}_{cross}.$ (14)
This pre-training strategy enables CCL-SLR to align embedding spaces across modalities, performing instance discrimination for consistency representation learning.
3.4 Model Details
During the pre-training phase, each projection head is a 2-layer MLP which projects the encoder outputs into a 128-dimensional embedding. During the fine-tuning phase, we employ the query encoders as feature extractors and replace the projection head with a fully connected layer for prediction. In this phase, both branches are optimized via cross-entropy loss. The final prediction is obtained by summing the predicted logits from the RGB and pose branches.
4 Experiments
Method Modality WLASL2000 WLASL300 MSASL1000 MSASL200
Pose RGB Per-instance Per-class Per-instance Per-class Per-instance Per-class Per-instance Per-class
Top-1 Top-5 Top-1 Top-5 Top-1 Top-5 Top-1 Top-5 Top-1 Top-5 Top-1 Top-5 Top-1 Top-5 Top-1 Top-5
ST-GCN [40] 34.40 66.57 32.53 65.45 44.46 73.05 45.29 73.16 36.03 59.92 32.32 57.15 52.91 76.67 54.20 77.62
I3D [1] 32.48 57.31 - - 56.14 79.94 - - - - 57.69 81.08 - - 81.97 93.79
TCK [28] - - - - 68.56 89.52 68.75 89.41 - - - - 80.31 91.82 81.14 92.24
StepNet [36] 56.89 88.64 54.54 87.97 74.70 91.02 75.32 91.17 - - - - - - - -
StepNet [36] 61.17 91.94 58.43 91.43 - - - - - - - - - - - -
HMA [14] 51.39 86.34 48.75 85.74 - - - - 69.39 87.42 66.54 86.56 85.21 94.41 86.09 94.42
BEST [46] 54.59 88.08 52.12 87.28 75.60 92.81 76.12 93.07 71.21 88.85 68.24 87.98 86.83 95.66 87.45 95.72
SignBERT [13] 54.69 87.49 52.08 86.93 74.40 91.32 75.27 91.72 71.24 89.12 67.96 88.40 86.98 96.39 87.62 96.43
SignBERT+ [16] 55.59 89.37 53.33 88.82 78.44 94.31 79.12 94.43 73.71 90.12 70.77 89.30 88.08 96.47 88.62 96.47
SAM($5^{\dagger}$[20] 58.73 91.46 55.93 90.94 - - - - - - - - - - - -
SAM-v2($5^{\dagger}$[21] 59.39 91.48 56.63 90.89 - - - - - - - - - - - -
NLA-SLR [55] 61.05 91.45 58.05 90.70 86.23 97.60 86.67 97.81 72.56 89.12 69.86 88.48 88.74 96.17 89.23 96.38
NLA-SLR($3^{\dagger}$[55] 61.26 91.77 58.31 90.91 86.98 97.60 87.33 97.81 73.80 89.65 70.95 89.07 89.48 96.69 89.86 96.93
CCL-SLR (Ours) 62.20 92.39 58.80 91.85 86.28 97.63 86.81 97.87 77.71 92.28 76.27 92.30 90.21 96.47 90.86 96.65
[…]
Method Modality NMFs-CSL SLR500
Pose RGB Top-1 Top-5 Top-1
ST-GCN [40] 59.9 86.8 90.0
I3D [1] 64.4 88.0 -
GLE-Net [15] 69.0 88.1 96.8
StepNet [36] 77.2 92.5 -
StepNet [36] 83.6 97.0 -
HMA [14] 75.6 95.3 95.9
SignBERT [13] 78.4 97.3 97.6
SignBERT+ [16] - - 97.8
BEST [46] 79.2 97.1 97.7
NLA-SLR [55] 83.4 98.3 -
NLA-SLR(3[55] 83.7 98.5 -
CCL-SLR (Ours) 84.4 99.3 97.8
[…]
4.1 Datasets and Evaluation Metrics
We evaluate our method on four public sign language datasets: MSASL [24], WLASL [26], NMFs-CSL [15], and SLR500 [18], utilizing the entire training sets of all four datasets to pre-train our framework.
MSASL is an American Sign Language (ASL) dataset, consisting of 25,513 samples with a vocabulary size of 1,000. It includes 16,054, 5,287, and 4,172 samples in the training, development (dev), and test set, respectively. Additionally, it provides a subset containing only the top 200 most frequent glosses. WLASL is another widely used ASL dataset with a vocabulary size of 2,000. It includes 14,289, 3,916, and 2,878 samples in the training, dev, and test sets, respectively. Similar to MSASL, WLASL also provides a subset of 300 frequent glosses.
NMFs-CSL is a challenging Chinese Sign Language (CSL) dataset that incorporates a wide range of fine-grained non-manual features (NMFs). With a vocabulary size of 1,067, it contains 25,608 training samples and 6,402 test samples. SLR500 is another large-scale CSL dataset covering 500 common words. It contains a total of 125,000 samples, with 90,000 samples for training and 35,000 samples for testing. Following [13, 55], we report the average accuracy over instances and classes, respectively. Since NMFs-CSL and SLR500 exhibit a balanced distribution with an equal number of samples per class, we only report per-instance accuracy.
4.2 Implementation Details
The RGB branch applies spatial augmentations including random cropping, horizontal flipping, and color jittering, while the pose branch implements random rotation, scaling, joint masking, and flipping. Both branches share temporal augmentation through random temporal cropping. More details about data preprocessing, the pose encoder and data augmentation are provided in the Appendix.
The temperature hyperparameters $\tau$ and $\tau_{spm}$ are set to $0.07$ and $0.02$, respectively. The hyperparameters of MPM $p$ and $\alpha$ are set to $0.5$ and $0.2$, respectively. During the pre-training phase, CCL-SLR is trained for 140 epochs with a learning rate of 0.01 and batch size 64, with the learning rate reduced by a factor of 10 at epoch 100. During the fine-tuning phase, the model is trained for 40 epochs with a learning rate of 0.05 and a batch size of 128, with the learning rate reduced by a factor of 10 at epochs 25 and 35. A Stochastic Gradient Descent (SGD) optimizer with a momentum of 0.9 is utilized throughout the training process. All experiments are conducted on 8$\times$ NVIDIA RTX 3090 GPUs.
4.3 Comparison with State-of-the-art Methods
MSASL. Table 1 presents an extensive comparison of our method against other approaches on two MSASL sub-splits. Benefiting from the consistency modeling between dual modalities during pre-training, our method surpasses the previously best-performing NLA-SLR [55], which utilized semantic information within sign glosses, by 5.15%/1.53% on the 1,000/200 sub-splits, respectively, in terms of per-instance top-1 accuracy.
WLASL. In Table 1, we also illustrates the results on the WLASL dataset. our CCL-SLR achieves a 2.81% improvement in the accuracy of the top-1 per instance compared to SAM-v2 [21], a multi-modal ensemble approach using RGB, pose, optical flow, and depth, demonstrating that alignment of the cross-modal distribution during pre-training enhances the generalization capability.
NMFs-CSL. The comparison on the NMFs-CSL dataset is shown in Table 2. Unlike previous pre-training methods such as SignBERT[13] or BEST[46], which simply fuse RGB predictions in downstream tasks and lack consideration of modeling cross-modal consistency during pre-training. Our method achieves superior performance with 6.0% and 5.2% improvements over SignBERT and BEST on the NMFs-CSL dataset, respectively.
SLR500. Lastly, Table 2 also presents the performance comparison on the SLR500 dataset. On this sufficiently large dataset, self-supervised learning methods such as SignBERT+[16] and BEST[46] achieve more outstanding performance. Similarly, our cross-modal pre-training approach effectively integrates consistency cues from both modalities, achieving a top-1 accuracy of 97.8%.
$\mathcal{L}^{m}_{spm}$ $\mathcal{L}_{cross}$ MPM Per-instance Per-class
Top-1 Top-5 Top-1 Top-5
70.04 88.43 67.73 86.91
71.20 88.93 68.96 88.14
76.27 91.78 74.26 91.57
77.71 92.28 76.27 92.30
Table 3: Ablation studies on the major components of CCL-SLR.
Setting Per-instance Per-class
Top-1 Top-5 Top-1 Top-5
w/o $\mathcal{L}_{cross}$ 72.69 90.23 79.81 89.46
w/ $\mathcal{L}_{cross}$ 77.71 92.28 76.27 92.30
Table 4: Ablation studies on cross-modal contrastive learning loss.
4.4 Ablation Study
In this section, we conduct several ablation studies to verify the effectiveness of key components in our proposed framework on the MSASL [48] dataset.
Key Components of CCL-SLR. As shown in Table 3, we conduct ablation studies on three key components of CCL-SLR: Semantic Positive Mining ($\mathcal{L}^{m}_{spm}$), cross-modal contrastive learning ($\mathcal{L}_{cross}$), and Motion-Preserving Masking (MPM). Initially, with a baseline comprising two independent single-modal learning branches, the incorporation of SPM improves top-1 accuracy by 0.85%, which is attributed to semantic positive pairs enforcing cross-modal consistency between sample spaces. The implementation of $\mathcal{L}_{cross}$ further boosts the performance significantly from 74.66% to 76.27%, demonstrating its effectiveness in aligning multi-modal representation during pre-training. Ultimately, by suppressing non-semantic features in RGB videos, the integration of MPM achieves the optimal performance with 77.71% top-1 accuracy.
Cross-Modal Contrastive Learning Loss. As shown in Table 4, we first investigate the impact of the cross-modal contrastive learning loss $\mathcal{L}_{cross}$. After removing $\mathcal{L}_{cross}$ from CCL-SLR, the top-1 accuracy significantly drops by about 5%. The reason is that $\mathcal{L}_{cross}$ plays a crucial role in our framework by explicitly and effectively aligning the feature spaces between the two modalities, which implicitly helps the model extract cross-modal consistency cues. Additionally, Fig. 4 demonstrates the heatmaps generated by fine-tuning RGB encoder from pre-training with and without $\mathcal{L}_{cross}$. Especially, performing cross-modal contrastive learning during pretraining allows the RGB branch to learn semantic information from the pose modality. This significantly enhances the model’s attention to non-manual features (yellow and red areas) such as facial expressions and body movements.
Figure 4: Visualizations of heatmaps by Grad-CAM [35]. Top: raw frames; Middle: heatmaps of the fine-tuning results of the RGB encoder from pre-training without $\mathcal{L}_{cross}$; Bottom: heatmaps of the fine-tuning results of the RGB encoder from pre-training with $\mathcal{L}_{cross}$. All samples are from the NMFs-CSL [15] dataset.
Pre-Training Data Scale. In Fig. 5, we examine the impact of the pre-training data scale on the performance of our proposed framework. The experimental results reveal that the performance exhibits a steady improvement as the proportion of the pre-training data scale increases, demonstrating that our model can effectively leverage large-scale unlabeled multi-modal sign language data to enhance its representation capability, highlighting the scalability of our approach to large-scale sign language data.
Figure 5: Hyperparameter selection on CCL-SLR performance. We systematically investigate the impact of pre-training data scale, mask probability $\alpha$ in MPM, and number of pseudo-positive labels $k$ in SPM.
Motion-Preserving Masking. The influence of MPM is explored by sequentially changing the mask probability parameter $\alpha$ within $[0.0,1.0]$. As shown in Fig.5, the performance reaches its peak at $\alpha=0.2$, indicating that moderate masking effectively mitigates static bias while preserving essential dynamic information. However, an excessively high mask ratio can lead to performance degradation. Visualization results of MPM are provided in the Appendix.
Semantic Positive Mining. We investigate the sensitivity of hyperparameter $k$ to our method, which denotes the number of mined semantic positives per modality. As illustrated in Fig. 5, experimental results demonstrate that either extremely small or large values of $k$ lead to performance degradation. An overly large $k$ will introduce noise during contrastive learning, negatively impacting the representation capability of the model. Based on these empirical findings, we set $k=10$ as the default value.
Setting Per-instance Per-class
Top-1 Top-5 Top-1 Top-5
$\mathcal{M}_{f}$ only 76.94 92.01 75.86 92.02
$\mathcal{M}_{g}$ and $\mathcal{M}_{f}$ 77.71 92.28 76.27 92.30
Table 5: Ablation studies on cross-modal contrastive learning loss.
Setting Fine-tuning Per-instance Per-class
Pose RGB Top-1 Top-5 Top-1 Top-5
Baseline 62.96 83.13 60.48 82.60
60.25 82.78 57.16 80.35
70.04 88.43 67.73 86.91
CCL-SLR 71.98 89.93 70.04 89.41
68.14 86.58 65.43 86.21
77.71 92.28 76.27 92.30
Table 6: Ablation studies on different pre-training modalities.
Figure 6: t-SNE visualizations of feature embeddings. We sample 30 sign words from the SLR500 dataset and visualize the feature embeddings extracted by CCL-SLR and its baseline, respectively.
Memory Banks. Unlike the vanilla MoCo v2 [3] framework, our pipeline maintains two memory banks ($\mathcal{M}_{g}$ and $\mathcal{M}_{f}$) in our pipeline: one for pseudo-labeling and the other for contrastive learning. Intuitively, maintaining a unique memory bank is structurally more concise. However, according to previous work [9], features before projection are more generally applicable, while features after projection are specialized for the instance discrimination task in contrastive learning. As shown in the Table 5, our experimental results further confirm that directly using $\mathcal{M}_{f}$ for pseudo-labeling leads to a slight performance degradation.
Cross-Modal Consistency Pre-training. As illustrated in Table 6, our framework is pre-trained under two different settings: baseline and CCL-SLR. The baseline conducts pre-training on two single-modal branches without key components of CCL-SLR. Compared with baseline, our CCL-SLR significantly improves the ISLR performance on the MSASL dataset for both single-modal and fusion results, which validates that our method effectively learns consistency representation across modalities. To further validate the effectiveness of our cross-modal consistency learning, t-SNE visualizations of the feature embeddings learned from pre-training are shown in Fig. 6. For both modalities, the representations learned by our CCL-SLR demonstrate more compact clustering patterns compared to the baseline method, indicating the superior representation learning capability of CCL-SLR.
5 Conclusion
In this paper, we investigate the challenge of sign language recognition through cross-modal consistency learning. We propose CCL-SLR, a novel self-supervised multi-modal pre-training framework that enhances sign representation by leveraging consistency information across RGB and pose modalities. Additionally, our framework incorporates two key innovations: MPM, which enables motion-aware regions extraction from RGB videos to tackle the problem of redundant features, and SPM, which effectively addresses the semantic collision problem in contrastive learning. We conduct comprehensive evaluations on four widely used ISLR datasets, where CCL-SLR consistently achieves impressive performance, thereby validating the effectiveness of our proposed approach.
Supplementary Material
A Framework Implementation
In this section, we describe the details of the pose data extraction, the encoder architecture in the pose branch and data preprocessing.
A.1 Pose Keypoint Extraction.
We employ RTMPose-x [22] from MMPose to extract 133 whole-body keypoints, which provide detailed spatial information for subsequent pose analysis. The visualization of whole-body keypoints are shown in Figure 7. We select 75 keypoints and divide the keypoints into several sub-pose (left hand, right hand, face, mouth, and body). We select the indices for the left hand ({92-112}), right hand ({113-133}), face ({24-41, 54}), mouth ({84-92}) and body ({0,5,7,9,6,8,10}) to represent each group (denoted as $\mathcal{G}_{r},r\in\{LH,RH,F,M,B\}$).
A.2 Encoder of Pose Branch.
We utilize a GCN+Transformer architecture [50] as the pose branch encoder. Specifically, four ST-GCN modules [41] are employed to extract group-specific features $\mathbf{v}_{r}$ from pose group $\mathcal{G}_{r}$, where the left and right hands share the same module for feature extraction. The extracted 512 dimension features are then concatenated into manual features $\mathbf{v}_{man}$ and non-manual features $\mathbf{v}_{non}$:
$\mathbf{v}_{man}$ $=\text{Concat}([\mathbf{v}_{LH},\mathbf{v}_{RH},\mathbf{v}_{B}]),$ (15)
$\mathbf{v}_{non}$ $=\text{Concat}([\mathbf{v}_{F},\mathbf{v}_{M}]),$ (16)
where $\mathbf{v}_{LH}$, $\mathbf{v}_{RH}$, $\mathbf{v}_{B}$, $\mathbf{v}_{F}$, and $\mathbf{v}_{M}$ represent features from left hand, right hand, body, face, and mouth, respectively. To model temporal relationships, two separate 3-block 8-head Transformers ($\mathcal{T}_{man}$ and $\mathcal{T}_{non}$) are leveraged to process these features and generate the final pose embedding:
$\mathbf{v}=\text{Concat}([\mathcal{T}_{man}(\mathbf{v}_{man}),\mathcal{T}_{non}(\mathbf{v}_{non})]),$ (17)
where final pose embedding $\mathbf{v}$ is 1024 dimensions.
A.3 Data Preprocessing.
For data preprocessing, we select aforementioned 75 keypoints for pose input per frame and resize all RGB frames to $224\times 224$. For computational efficiency, we uniformly sample $T=32$ frames from each sequence.
A.4 Data Augmentation
For the query and key samples, we ensure consistency in hyperparameters during data augmentation while performing the process independently. For random temporal cropping, we first apply continuous temporal cropping to the input sequence, randomly extracting a clip from the interval $[lT,T]$ frames. Subsequently, we uniformly sample a fixed number of $K$ frames from the cropped interval. In our approach, we set $l=0.1$ and $K$ = 64.
Figure 7: The visualization of the whole-body 133 keypoints from  [23].
B Visualization of Motion-Preserving Masking
As shown in Fig. 8 and Fig. 9, the generated motion-preserving videos effectively suppress static regions while highlighting motion areas. Furthermore, since motion-preserving videos significantly alter the pixel distribution of the original videos, we further generate binary motion-preserving mask sequences from the motion-preserving videos. These mask sequence are applied to mask original videos, explicitly mitigating static information redundancy in RGB modality. The visualizations indicate that MPM preserves semantically informative regions such as hand and facial features, while effectively suppressing semantically irrelevant areas including clothing textures and background elements, enhancing the cross-modal feature alignment between pose and RGB representations.
Figure 8: The visualizations of motion-preserving masking.
Figure 9: The visualizations of motion-preserving masking.